Given a current of 1.3[A]; a coil radius of 0.16[m
Answer #1


5) I =1.3 A, r = 0.16 m, N =130

B = Nuo*i/2a

B = 130*4*3.14*10^-7*1.3/(2*0.16)

B = 0.66 mT

6) B =0.00095 T , V/d^2 = 19629 V/m^2

mv^2/r = qvB

q/m = v/rB ..(1)

From conservation of energy

K =U

1/2 mv^2 = qV

v = (2qV/m)^0.5

From(1)

q/m = (2qV/m)^0.5 /rB

q/m = 2V/r^2B^2

e/m = 2*19629/(0.00095^2)

e/m = 0.435*10^11 C/kg


7) e/m = 8 B^-28*V*d^-2

correct option is (b)